Statistics

The U.S. Bureau of Labor Statistics released hourly wage figures for various countries for workers in the manufacturing sector. The hourly wage was $30.67 for Switzerland, $20.20 for Japan, and $23.82 for the U.S. Assume that in all three countries, the standard deviation of hourly labor rates is $3.00.
Answer #1

a)

for normal distribution z score =(X-μ)/σx


here mean=       μ=30.67
std deviation   =σ=3.000
sample size       =n=41
probability =P(30<X<31)=P(-1.43<Z<0.7)=0.758-0.0764=0.6816

b)

here mean=       μ=20.2
std deviation   =σ=3
sample size       =n=32
std error=σ=σ/√n=0.5303
probability =P(X>21)=P(Z>1.51)=1-P(Z<1.51)=1-0.9345=0.0655

c)

probability =P(X<22.75)=P(Z<-2.45)=0.0071

d)

probability =P(X<22.9)=P(Z<-2.12)=0.0170


answered by: Keynes